At Trapeze Corporation, direct materials are added at the beginning of the process and conversions costs are uniformly applied (60% complete at the end of the month).

WIP beginning 0 units, $0
Units started 122,500 units
Units completed 116,700 units
Costs of materials added $382,720
Costs of conversion added $275,790

a. Calculate equivalent units for direct materials and conversion costs.

b. Calculate the cost per unit of direct materials and cost per unit of conversion costs.

c. What is the total amount of costs transferred to finished goods inventory?

d. What is the total amount of WIP ending balance in dollars?
Answer #1

a) Equivalent unit of material = 116700+(5800*100%) = 122500

Equivalent unit of conversion = 116700+(5800*60%) = 120180

b) Cost per equivalent unit of material = 382720/122500 = 3.12

Cost per equivalent unit of conversion = 275790/120180 = 2.29

c) Cost transferred to finished goods = (3.12+2.29)*116700 = $631347

d) Ending WIP = (5800*3.12+3480*2.29) = 26065